Jerome is 4 times old as Rugem. Ten years from now, he will be twice as old as Rugem will be. How old is Jerome now?
Answer by gsmani_iyer(201)

Let the age of Rugem now be = x
So the age of Jerome now be = 4x ....... 4 times the age of Rugem
After 10 years, the ages of
Rugem will be = x + 10
Jerome will be = 2(x + 10) ..........Twice the age of Rugem
As per the first equation,
the age of Jerome is = 4x +10
So, 2(x +10) = 4x + 10
= 2x + 20 = 4x + 10
= 20 - 10 = 4x - 2x
= 2x = 10
= x = 10/2 = 5
So the present age of Regem = 5 and
the present age of Jerome = 5*4 = 20 years Answer

I hope this is clear to you. In such problems, randomly you can verify
the present and the age after 10 years applying the values found out.


GSM